Suppose1,006 students attend middle school where ms.yadav teaches. Use your answer from part (a) to estimate the number of students who are wearing boots. Explain.


1
Expert's answer
2021-12-09T03:02:14-0500

(a)

There are 30 students in Ms. Yadav's math class. Today in math class, 16 students are wearing sneakers, 10 are wearing boots, and 4 are wearing other kinds of shoes.

What fraction of Ms. Yadav's students are wearing boots?


"\dfrac{10}{30}=\dfrac{1}{3}"

(b)

Suppose1,006 students attend middle school where ms.yadav teaches. Use your answer from part (a) to estimate the number of students who are wearing boots.

If we assume that the ratio of students in Ms. Yadav's class wearing boots is the same as the ratio of students in entire school wearing boots, then an estimate would be


"\dfrac{1}{3}(1006)=335"

335 students are wearing boots.